# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

### v3

@@ 4.3 Generalization Experiments
Besides the ability to follow soft constraints, we also assess the model’s general instruction following abilities on AlpacaEval2.0 Zhao et al. (2024a) and MT-Bench Zheng et al. (2023). We first perform SFT on the model, followed by DPO+Judger+CL. Specifically, we use precomputed outputs of GPT-4 Turbo on AlpacaEval as reference outputs and GPT-4o as evaluators. As shown in the Tab. 4.3, our method improves the model’s general instruction following ability on both banchmarks.
@@ 4.4 Ablation Studies
Curriculum Setting FollowBench (HSR) IFEval
L1 - L3 L4 - L5 Avg Avg
BASE 56.3 44.9 51.7 74.3
{C1}{C2}{C3}{C4}{C5} 57.4 37.9 49.6 76.7
{C1,C2,C3}{C4,C5} 59.9 46.8 54.6 78.5
Table 6: Results of different curriculum setting.
@@
We also explore the impact of curriculum setting. As shown in Tab. 6, combining Curriculum 1 to Curriculum 3 into an easy curriculum and Curriculum 4 to Curriculum 5 into a difficult curriculum significantly improves model’s performance on both benchmarks compared to five separate curricula, especially on FollowBench. If the model is trained on five separate curricula, there is no clear boundary between task complexities. The model struggles to generalize from easier tasks to more difficult ones. Grouping similar tasks together with a clear progression in difficulty helps the model better focus on mastering each level of complexity, preventing overfitting to individual tasks.
4.5 Analysis
Figure 4: Results across various constraint categories.
4.5.1 Category Analysis
In this section, we analyze the model’s performance across different types of constraints. Specifically, we compare the performance of LLaMA3-8B-Instruct with the BASE and DPO+Judger+CL method on FollowBench. As shown in Fig. 4, our method significantly improves the model’s performance across different types of constraints. The most notable improvement is observed in the Mixed category , which is defined as a composition of multiple constraint categories to simulate complex real-world scenarios. Our method enhances the model’s performance by 10.6% in this category, suggesting a notable enhancement in the model’s ability to handle complex constraints in real-world scenarios. For categories that contain only soft constraints, our method improves the model’s performance by 6% in the Style category. However, in the category of situation constraints, the performance of the DPO+Judger+CL method is worse than that of BASE. The reason is that our constructed situation constraints focus on soft constraints, while in FollowBench, the situation constraints include both soft and hard constraints. Our method struggles with the hard constraints in the situation category, such as complex situation reasoning, leading to the decrease in performance.
4.5.2 The Role of Progressive Construction
@@ 4.5.3 The Role of Judger
In this section, we investigate the role of Judger in obtaining high-quality outputs. Specifically, we randomly select 100 seed instructions and rank the outputs at each step of progressive construction . We evaluate the rankings in three scenarios: (1) w/o Judger: directly use GPT-4o to generate outputs without reordering, (2) w/ Judger: use Judger to reorder the outputs, and (3) rankings annotated by human experts, which serve as the reference standard for comparison.
To assess the similarity between these rankings, we employ two metrics. The first is the Kendall Tau Coefficient Kendall (1938),which measures the correlation between two rankings by assessing the agreement in the order of paired items. Another metric is position consistency, which quantifies the proportion of elements that occupy the same relative positions. As shown in Tab. 8, the rankings adjusted by the Judger exhibit greater alignment with human-annotated rankings. This suggests that Judger enhances the consistency of outputs with human judgments, thereby improving their quality.
@@ 5 Conclusion
In this paper, we systematically study how to improve LLMs’ ability to follow soft constraints. Initially, we design a pipeline to automate the construction of datasets with high-quality outputs for soft constraint following. Based on the pipeline, we introduce a method utilizing positive and negative samples generated during the pipeline. Moreover, we propose a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. The experiment results show that our methods enhance models’ ability to follow soft constraints effectively.
@@ A.2 Details of Judger Reordering
Judger Ranking Case
Instruction Please write a brief description of quantum physics, ensuring that the terms "wave function" and "superposition state" are included. Avoid using overly complex formulas. Keep the response under 100 words.Ensure that this description is easy to understand.
Model Responses Output (a) : Quantum physics studies the behavior of tiny particles. The *wave function* describes a particle’s state, while *superposition state* means a particle can exist in multiple states at once.
Output (b) : Quantum physics studies the behavior of tiny particles. In the quantum world, particles don’t have fixed positions or states; they can be in many states at once, and only take a definite state when measured.
Judger Output [[A]]. From the perspective of constraint adherence, Output (a) is better. It clearly includes both required terms, "wave function" and "superposition state," and explains them in a simple and concise way. The response stays under 100 words and avoids complex formulas, making it easy to understand.Output (b), on the other hand, doesn’t mention "wave function" and its explanation of "many states" is less clear, making it less aligned with the instruction.
Table 9: The case for Judger ranking outputs.
@@ A.3 Implementation Details
We train LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and LLaMA2-13B-Chat-HF using LLaMA-Factory Zheng et al. (2024) on 4 NVIDIA A100 80GB GPUs, applying LoRA Hu et al. (2021) for efficient training. The lora target is set to all, with all models training for 3 epochs. The per device train batch size is set to 1, and gradient accumulation steps is set to 8. The warm-up ratio is set to 0.1. For SFT, LLaMA-3-8B-Instruct has a learning rate of 1.0e-4, Mistral-7B-Instruct-v0.3 uses 5.0e-7, and LLaMA2-13B-Chat-HF uses 1.0e-7. For DPO, the learning rate is 5.0e-6 with a beta value of 0.1. We apply cosine learning rate scheduler.

### v4

@@ 4.3 Generalization Experiments
Besides the ability to follow soft constraints, we also assess the model’s general instruction following abilities on AlpacaEval2.0 Zhao et al. (2024a) and MT-Bench Zheng et al. (2023). We first perform SFT on LLaMA3-8B-Instruct, followed by DPO+Judger+CL. Specifically, we use precomputed outputs of GPT-4 Turbo on AlpacaEval as reference outputs and GPT-4o as evaluators. As shown in the Tab. 4.3, our method improves the model’s general instruction following ability on both banchmarks.
@@ 4.5.1 The Role of Progressive Construction
Judger Ranking Case
Instruction Please write a description of quantum physics, including "wave function" and "superposition state", avoiding complex formulas, under 100 words. Ensure this description is easy to understand.
Responses (a) Quantum physics studies the behavior of tiny particles. The *wave function* describes a particle’s state, while *superposition state* means a particle can exist in multiple states at once.
(b) Quantum physics studies the behavior of tiny particles. In the quantum world, particles don’t have fixed positions or states; they can be in many states at once, and only take a definite state when measured.
Judger Output (a) is better. It includes required terms, explains them simply, and follows all constraints. Output (b) misses "wave function" and is less precise.
Table 10: The case for Judger ranking outputs.
@@ 4.5.2 The Role of Judger
In this section, we investigate the role of Judger in obtaining high-quality outputs. Specifically, we randomly select 100 seed instructions and rank the outputs at each step of progressive construction. We evaluate the rankings in three scenarios: (1) w/o Judger, (2) w/ Judger and (3) rankings annotated by human experts, which serve as the reference standards. As shown in Tab. 10, Output(a) represents the model’s initial response generated without the newly introduced constraint “Ensure that this description is easy to understand” while Output(b) shows the response produced after incorporating this additional constraint. In the w/o Judger setting, it assumes Output(b) to be superior by default, without performing reordering. To assess the similarity between the rankings, we employ two metrics. The first is the Kendall Tau Coefficient Kendall (1938), which measures the correlation between two rankings by assessing the agreement in the order of paired items. Formally, given two rankings $R_{1}$ and $R_{2}$ over the same set of $n$ items. Kendall Tau Coefficient is defined as
$\tau(R_{1},R_{2})=\frac{\displaystyle\sum_{1\leq i<j\leq n}\operatorname{sgn}\bigl{(}R_{1}(i)-R_{1}(j)\bigr{)}\,\operatorname{sgn}\bigl{(}R_{2}(i)-R_{2}(j)\bigr{)}}{\tfrac{1}{2}\,n(n-1)}$
The second metric is Position Consistency, which quantifies the proportion of elements that occupy the same relative positions in both rankings. If $\mathrm{pos}_{R_{1}}(i)$ and $\mathrm{pos}_{R_{2}}(i)$ denote the positions of item $i$ in $R_{1}$ and $R_{2}$, respectively, then
$\mathrm{PC}(R_{1},R_{2})=\frac{1}{n}\,\bigl{|}\{\,i\mid\mathrm{pos}_{R_{1}}(i)=\mathrm{pos}_{R_{2}}(i)\}\bigr{|}$
As shown in Tab. 9, the rankings adjusted by the Judger exhibit greater alignment with human-annotated rankings. This suggests that Judger enhances the consistency of outputs with human judgments, thereby improving their quality.
@@ 5 Conclusion
In this paper, we systematically study how to improve LLMs’ ability to follow soft constraints. Initially, we design a pipeline to automate the construction of datasets with high-quality outputs for soft constraint following. Based on the pipeline, we introduce a reinforcement learning training method utilizing positive and negative samples generated during the pipeline. Moreover, we propose a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. The experiment results show that our methods enhance models’ ability to follow soft constraints effectively.
@@ A.3 Implementation Details
We train LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and LLaMA2-13B-Chat-HF using LLaMA-Factory Zheng et al. (2024) on 4 NVIDIA A100 80GB GPUs, applying LoRA Hu et al. (2021) for efficient training. The lora target is set to all, with all models training for 3 epochs. The per device train batch size is set to 1, and gradient accumulation steps is set to 8. The warm-up ratio is set to 0.1. For SFT, LLaMA-3-8B-Instruct has a learning rate of 1.0e-4, Mistral-7B-Instruct-v0.3 uses 5.0e-7, and LLaMA2-13B-Chat-HF uses 1.0e-7. For DPO, the learning rate is 5.0e-6 with a beta value of 0.1. We apply cosine learning rate scheduler. For the benchmark evaluation, we adopt the online serving inference approach using vLLM.